# Random Lévy Looptrees and Lévy Maps

### v2

@@ 2.3.   Invariance principles
    From every increasing sequence of integers, one can extract a subsequence along which the convergence in distribution
    $\Bigl{(}\frac{1}{\sqrt{2r_{n}}}d_{M^{n}}(v^{n}_{\lfloor ns\rfloor},v^{n}_{\lfloor nt\rfloor})\Bigr{)}_{s,t\in[0,1]}\enskip\mathop{\longrightarrow}^{(d)}_{n\to\infty}\enskip(D(s,t))_{s,t\in[0,1]},$
    holds for the uniform topology, where $D$ is a random continuous pseudo-distance which satisfies
    $D(U,\cdot)\enskip\mathop{=}^{(d)}\enskip Z^{\mathrm{ex}}-\min Z^{\mathrm{ex}},$
    where $U$ is sampled uniformly at random on $[0,1]$ and independently of the rest. Finally along this subsequence, we have
    $\frac{1}{\sqrt{2r_{n}}}M^{n}\enskip\mathop{\longrightarrow}^{(d)}_{n\to\infty}\enskip{}M=[0,1]/\{D=0\}$
    in the Gromov–Hausdorff–Prokhorov topology.
@@
    If $X^{\mathrm{ex}}$ has no jump, so it reduces to $\beta$ times the Brownian excursion, then these subsequential limits agree with $\sqrt{\beta/3}$ times the standard Brownian sphere.
@@
The first claim is a straightforward consequence of Theorem 2.1, see e.g. [Mar22b, Lemma 2.5]. The second claim is not immediate but is very standard since the original work of Le Gall [LG07], see also [Mar22b, Lemma 2.6] for a recent account of the argument is a similar context. As explained in Remark 2.3, this lemma requires an extra condition in order to remove the redundancies in the list $(v^{n}_{0},\dots,v^{n}_{n})$ of vertices, see [Mar22b, Equation 4]. As recalled there, the latter is provided by [Mar22a, Lemma 2.2] with $A=\{-1\}$. The last claim is a consequence of the identification of the Brownian sphere and the argument can be found in [LG13], see precisely the proof of Equation 59 there. ∎
@@
In the rest of this paper, we call a Lévy map any subsequential limit $M$ in the previous theorem. We equip it with the metric $d_{M}$, the image of $D$ by the canonical projection, and the probability $p_{M}$, given by the push-forward of the Lebesgue measure on $[0,1]$. Similarly to looptrees in Remark 2.2, Lévy maps are geodesic spaces.
@@ 3. A spinal decomposition and some volume bounds
We then turn to maps. Recall that we let $(M,d_{M},p_{M})$ denote a subsequential limit of discrete random maps, related to a pair $(X^{\mathrm{ex}},Z^{\mathrm{ex}})$, from Theorem 2.1 and Theorem 2.4. In addition, thanks to Skorokhod’s representation theorem, we assume that all convergences in these theorems hold in the almost sure sense. Notice the additional square inside $\psi$ in the next statement compared to the first one.
@@
In order to prove these results, we shall first describe Section 3.1 the path from the root to a uniformly random point in the Lévy looptree, as well as the labels on it. Then Proposition 3.1 is proved in Section 3.2, whereas Proposition 3.2 is proved in Section 3.3, relying on a technical estimates on Brownian paths stated in Section 3.4 and proved in Section 3.5.
@@ 3.1.   A spinal decomposition
It is straightforward to check that these properties hold in each case on the right of the claim, let us prove the direct implication when $Y=X^{\mathrm{ex}}$ by relying on some properties of such a path.
@@
Consider $X$ the unconditioned Lévy process. First almost surely if $\Delta X_{r}>0$, then for every $\varepsilon>0$, we have both $\inf_{[r-\varepsilon,r]}X<X_{r-}$ and $\inf_{[r,r+\varepsilon]}X<X_{r}$. The latter follows from the Markov property and the fact that 00 is regular for $(-\infty,0)$ and the former by invariance under time reversal and the fact that 00 is regular for $(0,\infty)$, see Theorem VII.1 and Corollary VII.5 in [Ber96]. Consequently almost surely if $X$ achieves a local minimum, then it does not jump at this time. Next, the strong Markov property also entails that the local minima are unique. Finally, by [Ber96, Proposition III.2], we know that almost surely every weak record of $X$ obtained by a jump is actually a strict record. Applied to the dual process at a time just after $t$, it shows that almost surely if $X$ realises a local minimum at time $t$ and if $r=\sup\{u<t\colon X_{u}<X_{t}\}$ has $\Delta X_{r}>0$, then $X_{r-}<X_{t}<X_{r}$.
All these properties transfer to the bridge $X^{\mathrm{br}}$ by absolute continuity and then to the excursion $X^{\mathrm{ex}}$ by the Vervaat transform. Therefore on an event of probability one, the four conditions above simplify as:
  1. (i)
    For every $r$ such that $s\wedge t\prec r\prec s$, it holds $s=\inf\{u>r\colon X^{\mathrm{ex}}_{u}=X^{\mathrm{ex}}_{r-}\}$;
  2. (ii)
    For every $r$ such that $s\wedge t\prec r\prec t$, it holds $t=\inf\{u>r\colon X^{\mathrm{ex}}_{u}=X^{\mathrm{ex}}_{r-}\}$;
  3. (iii)
    It holds $\inf_{[s\wedge t,s]}X^{\mathrm{ex}}=\inf_{[s\wedge t,t]}X^{\mathrm{ex}}$;
  4. (iv)
    Finally $C_{s}=C_{t}=\inf_{[s,t]}C$.
@@
If $s\not\prec t$, then almost surely there exists $u\in(s,t)$ such that $X^{\mathrm{ex}}_{u}=\inf_{[s,t]}X^{\mathrm{ex}}<X^{\mathrm{ex}}_{s-}$. Let $r=s\wedge t$. Since $X^{\mathrm{ex}}$ realises a local minimum at time $u$, then almost surely we must have $\Delta X^{\mathrm{ex}}_{r}>0$ and further $X^{\mathrm{ex}}_{r-}<X^{\mathrm{ex}}_{u}<X^{\mathrm{ex}}_{s}<X^{\mathrm{ex}}_{r}$, which contradicts the third condition above. Therefore almost surely we have $s\prec t$, namely $X^{\mathrm{ex}}_{s-}\leq\inf_{[s,t]}X^{\mathrm{ex}}$. If $\Delta X^{\mathrm{ex}}_{s}>0$ then indeed the second condition implies that $t=\inf\{u>r\colon X^{\mathrm{ex}}_{u}=X^{\mathrm{ex}}_{r-}\}$. Otherwise if $\Delta X^{\mathrm{ex}}_{s}=0$, then there is no $r\in(s,t)$ such that $X^{\mathrm{ex}}$ achieves a local minimum at time $r$ with $X^{\mathrm{ex}}_{r}=X^{\mathrm{ex}}_{s}$, nor $X^{\mathrm{ex}}_{s}=X^{\mathrm{ex}}_{r-}<X^{\mathrm{ex}}_{r}$. Therefore if $X^{\mathrm{ex}}_{t-}>X^{\mathrm{ex}}_{s}$, and since $d_{C}(s,t)=0$, then there exists $r\in(s,t)$ such that $\Delta X^{\mathrm{ex}}_{r}>0$ and $X^{\mathrm{ex}}_{s}<X^{\mathrm{ex}}_{r-}$ and $X^{\mathrm{ex}}_{t-}<X^{\mathrm{ex}}_{r}$, which contributes to the second term in (2.6). We conclude that in this case we must have $X^{\mathrm{ex}}_{t-}=X^{\mathrm{ex}}_{s}$. Furthermore, the time $t$ cannot be a jump time, for otherwise $X^{\mathrm{ex}}$ would reach smaller values just before, hence $X^{\mathrm{ex}}_{t}=X^{\mathrm{ex}}_{s}=\inf_{[s,t]}X^{\mathrm{ex}}$. Finally there can only be at most two such times $t$, if $\inf\{t>s\colon X^{\mathrm{ex}}_{t}=X^{\mathrm{ex}}_{s}\}$ is a time of a local minimum since then it is unique at this height. ∎
@@
Let us now describe the ‘chain of loops’ from the root 00 to a given point $u\in[0,1]$ in the looptree, see Figure 7 for a graphical representation. First define the following set of pairs of times:
$A_{u}=\{(a,b)\in[0,u]\times[u,1]\colon\text{either }X^{\mathrm{ex}}_{a}=X^{\mathrm{ex}}_{b}=\min_{[a,b]}X^{\mathrm{ex}}\text{ or }b=\inf\{t>a\colon X^{\mathrm{ex}}_{t}=X^{\mathrm{ex}}_{a-}\}\}.$ (3.1)
@@
By Lemma 3.3, pairs $(a,b)\in A_{u}$ are identified in the looptree. These points intuitively correspond to the ancestors of $u$ in the tree. Define then the set of pairs of points that belong to the same cycle on the geodesic from $u$ to the root, with one point on the left and one on the right. Formally, set:
$B_{u}=\{(a,b)\in[0,u]\times[u,1]\colon\exists s<a\text{ such that }X^{\mathrm{ex}}_{s}>X^{\mathrm{ex}}_{a-}=\min_{[s,a-]}X^{\mathrm{ex}}>X^{\mathrm{ex}}_{b-}=\min_{[s,b-]}X^{\mathrm{ex}}>X^{\mathrm{ex}}_{s-}\}.$ (3.2)
@@
Pairs in $(a,b)\in B_{u}$ are not identified in the looptree (still by Lemma 3.3). Instead, the time $s$ in (3.2) together with $t=\inf\{v>s\colon X^{\mathrm{ex}}_{v}=X^{\mathrm{ex}}_{s-}\}>u$ belong to $A_{u}$, so they correspond to an ancestor of $u$. This ancestor has a large offspring number $\Delta X^{\mathrm{ex}}_{s}$ which creates a cycle in the looptree. One point on this cycle corresponds to the next ancestor of $u$ and splits the cycle into a left part, to which $a$ belongs, and a right part, to which $b$ belongs. One can easily prove that almost surely, for any $u\in[0,1]$ the set $A_{u}$ is closed and $A_{u}\cup B_{u}$ is the closure of $B_{u}$.
@@
    $u^{-}_{\varepsilon}=\inf\{s<u\colon X^{\mathrm{ex}}_{s-}=X^{\mathrm{ex}}_{s_{\varepsilon}-}=\min_{[s,u]}X^{\mathrm{ex}}\}\qquad\text{and}\qquad u^{+}_{\varepsilon}=\inf\{t>u\colon X^{\mathrm{ex}}_{t}<X^{\mathrm{ex}}_{s-}\},$ (3.3)

### v3

@@ 2.3.   Invariance principles
    From every increasing sequence of integers, one can extract a subsequence along which the convergence in distribution of continuous functions on $[0,1]^{2}$
    $\frac{1}{\sqrt{2r_{n}}}d_{M^{n}}\enskip\mathop{\longrightarrow}^{(d)}_{n\to\infty}\enskip D,$
    holds for the uniform topology, where $D$ is a random continuous pseudo-distance. Consequently, along this subsequence, the rescaled map $(2r_{n})^{-1/2}M^{n}$ converges in distribution to the quotient space $[0,1]/\{D=0\}$ for the Gromov–Hausdorff–Prokhorov topology.
@@
    Define for every $0\leq s\leq t\leq 1$:
    $D^{\circ}(s,t)=D^{\circ}(t,s)=Z^{\mathrm{ex}}_{s}+Z^{\mathrm{ex}}_{t}-2\max\Bigl{(}\inf_{[s,t]}Z^{\mathrm{ex}},\inf_{[0,s]\cup[t,1]}Z^{\mathrm{ex}}\Bigr{)}.$ (2.12)
    Then all the subsequential limits $D$ satisfy
    $D\leq D^{\circ}\qquad\text{almost surely}.$
  4. (iv)
    Sample $U$ uniformly at random on $[0,1]$ and independently of the rest. Then all the subsequential limits $D$ satisfy the identity in law:
    $(D(U,t))_{t\in[0,1]}\enskip\mathop{=}^{(d)}\enskip(Z^{\mathrm{ex}}_{t}-\min Z^{\mathrm{ex}})_{t\in[0,1]}.$
  5. (v)
    If $X^{\mathrm{ex}}$ has no jump, so it reduces to $\beta$ times the standard Brownian excursion, then we have the convergence without the need to extract a subsequence:
    $\frac{1}{\sqrt{2r_{n}}}d_{M^{n}}\enskip\mathop{\longrightarrow}^{(d)}_{n\to\infty}\enskip\sqrt{\beta/3}\cdot D^{\ast},$
    where $[0,1]/\{D^{\ast}=0\}$ is the standard Brownian sphere. Similarly, if $\psi(\lambda)=c\lambda^{\alpha}$ with $c>$ and $21<\alpha<$ 2, then the maps converge in distribution without extraction to $c^{1/(2\alpha)}$ times the standard $\alpha$-stable map.
@@
The first claim is a consequence of the convergence of $Z^{n}$ Theorem 2.1 since the distances to $v^{n}_{0}$ in the map are given by the labels on the looptree, recorded in $Z^{n}$, shifted by their overall minimum plus $1$, see e.g. the proof of Theorem 7.12 in [Mar22] for details. The second and third claims follow from an argument from the original work of Le Gall [LG07], see also [Mar24, Lemma 2.6] for a recent account of the argument in a similar context. In a few words, in the discrete models, recall that the map is constructed from the labelled looptree by linking each corner to its successor, defined as the first one with a smaller label when following the contour of the looptree. Then if we start from two corners, the two chains of successors eventually meet and the sum of the two lengths until this meeting point (minus $2$) is precisely the discrete analogue of $D^{\circ}$, say $D^{\circ}_{n}$, defined using $Z^{n}$, and this quantity upper bounds the distances in the map between the two vertices where we started. Since $D^{\circ}_{n}$ depends continuously on $Z^{n}$, then it converges in distribution after scaling to $D^{\circ}$, from which we easily derive both tightness of the distances in the map under the same rescaling due to the upper bound $d_{M^{n}}\leq D^{\circ}_{n}+2$, and $D\leq D^{\circ}$ after passing to the limit along any convergent subsequence.
The fourth claim is a re-rooting (or rather, re-pointing) invariance property: by the first two claims, the process on the right describes the distances to the distinguished vertex in the in the limit, whereas the one on the left describes the distances to an independent uniform random vertex. The identity thus follows from the fact that, in the discrete model, the distinguished vertex has the law of an independent uniform random vertex, which is due to the assumption (2.10).
The last claim in the Brownian case is a consequence of the identification of the Brownian sphere and the argument can be found in [LG13], see precisely the proof of Equation 59 there. It relies on the following ingredients: our claims 2, 3, and 4, the (nontrivial) fact that, if $U$ and $U^{\prime}$ are two independent uniform random times on $[0,1]$, then $Z^{\mathrm{ex}}_{U}-\min Z^{\mathrm{ex}}$ has the same law as $\sqrt{\beta/3}$ times the distances between $U$ and $U^{\prime}$ for the pseudo-distance $D^{\ast}$ describing the Brownian sphere, and finally, the fact that, a priori, for any subsequential limit $D$, we have $D(s,t)=0$ whenever $s$ and $t$ are identified in the Brownian tree coded by $X^{\mathrm{ex}}$. In a few words, we obtain from these ingredients the identity in law:
$D(U,V)\enskip\mathop{=}^{(d)}\enskip\sqrt{\beta/3}\cdot D^{\ast}(U,V),$
and in addition that $D\leq D^{\ast}$ almost surely so the above identity in law in fact holds almost surely so finally $D=\sqrt{\beta/3}\cdot D^{\ast}$ almost surely by a density argument. The same argument applies in the stable regime as discussed in [CMR25, Section 7.4.2] thanks again to the identification there of $D^{\ast}(U,U^{\prime})=Z^{\mathrm{ex}}_{U}-\min Z^{\mathrm{ex}}$ in law in this case, where in the last ingredient, we require that $D(s,t)=0$ whenever $s$ and $t$ are identified in the looptree tree coded by $X^{\mathrm{ex}}$. We will prove this last fact later in Lemma 4.8 in our general Lévy regime. ∎
The invariance by re-rooting argument used in the last claim in the Brownian and stable regimes applies in the general Lévy case and would similarly allow to obtain the convergence for all models whose Łukasiewicz paths converge to the same Lévy excursion, if one could prove the identity in law $D^{\ast}(U,U^{\prime})=Z^{\mathrm{ex}}_{U}-\min Z^{\mathrm{ex}}$, for example by proving the convergence of one specific discrete model of maps.
Remark 2.4.
The first part of the statement, characterising the distances to the distinguished vertex $v^{n}_{0}$, is not enough to prove that all subsequential limits agree. However, it does provide the convergence, without the need to extract a subsequence, of several quantities of interest. Let $V(M^{n})$ denote the set of vertices of $M^{n}$ and $\rho_{n}$ the origin of the root edge, then
$\frac{1}{\sqrt{2r_{n}}}d_{M^{n}}(v^{n}_{0},\rho_{n})\enskip\mathop{\longrightarrow}^{(d)}_{n\to\infty}\enskip-\min Z^{\mathrm{ex}}$
and
$\frac{1}{\sqrt{2r_{n}}}\max_{v\in V(M^{n})}d_{M^{n}}(v^{n}_{0},v)\enskip\mathop{\longrightarrow}^{(d)}_{n\to\infty}\enskip\max Z^{\mathrm{ex}}-\min Z^{\mathrm{ex}},$
and finally, for every continuous and bounded function $F$,
$\frac{1}{\#V(M^{n})}\sum_{v\in V(M^{n})}F\Bigl{(}\frac{1}{\sqrt{2r_{n}}}d_{M^{n}}(v^{n}_{0},v)\Bigr{)}\enskip\mathop{\longrightarrow}^{(d)}_{n\to\infty}\enskip\int_{0}^{1}F(Z^{\mathrm{ex}}_{t}-\min Z^{\mathrm{ex}})\mathop{}\!\mathrm{d}t.$
@@
In the rest of this paper, we call a Lévy map any subsequential limit $M=[0,1]/\{D=0\}$ in the previous theorem. We equip it with the metric $d_{M}$, the image of $D$ by the canonical projection, and the probability $p_{M}$, given by the push-forward of the Lebesgue measure on $[0,1]$. Similarly to looptrees in Remark 2.2, Lévy maps are geodesic spaces.
3. Limits of biconditioned planar maps and stable excursions with a drift
Let us present in this section applications of our general results in the case where the coding process is a spectrally positive stable Lévy process with a drift, defined just a few lines below, and recall how such objects appear as limits of multiconditioned discrete models introduced in [KM23].
for every $u\in\mathbb{R}$ and $t>0$, where we used the notation $\sigma(u)=\mathbf{1}_{\{u>0\}}-\mathbf{1}_{\{u<0\}}$ for the sign function to distinguish it from the sine function. The right-hand side satisfies the integrability condition (1.3), and thus it admits positive, bi-continuous, transition densities $p^{\alpha}_{t}$ given by inverse Fourier transform, from which one can define the conditional law $\mathop{}\!\mathbb{P}(\,\cdot\mid X^{\alpha}_{1}=\vartheta)$ for any $\vartheta\in\mathbb{R}$, characterised by (2.1). For any $\vartheta\in\mathbb{R}$, let us also denote by $X^{\alpha,\vartheta}_{t}=X^{\alpha}_{t}-\vartheta t$ the stable process with drift $-\vartheta$, which has transition densities $p^{\alpha}_{t}(\cdot+\vartheta t)$. One can also define bridges of this process and we let $X^{\alpha,\vartheta,\mathrm{br}}$ denote the version of $X^{\alpha,\vartheta}$ under the conditional law $\mathop{}\!\mathbb{P}(\,\cdot\mid X^{\alpha,\vartheta}_{1}=0)$. Finally let $X^{\alpha,\vartheta,\mathrm{ex}}$ denote the excursion obtained by applying the Vervaat transform (2.2) to $X^{\alpha,\vartheta,\mathrm{br}}$.
Recall from Section 2.1 the definition of Boltzmann random planar maps. From the bijection recalled there, Boltzmann maps with $n$ edges relate to certain (loop)trees and excursions with duration $n+1$ of downward skip-free random walks with step distribution $(\nu_{q}(k);k\geq-1)$ given by:
provided such a $Z_{q}>1$ exists to ensure that $\nu_{q}$ is indeed a probability. When $\nu_{q}$ is centred and has finite variance, the corresponding random walk excursion converges after scaling to the Brownian excursion, and accordingly, the Boltzmann maps conditioned to have many edges converge to the Brownian sphere, see e.g. [Mar22, Section 6.4] for a more general statement. However, when the probability $\nu_{q}$ is centred and has the asymptotic behaviour $\nu_{q}(k)\sim ck^{-1-\alpha}$ as $k\to\infty$ for some unimportant constant $c>0$ and an (important) index $\alpha\in(1,2)$, such as in the explicit example (1.1) in the introduction, then the excursion converges to that of $X^{\alpha}$. In this regime, Le Gall & Miermont [LGM11] proved Theorem 2.3 (i) and (ii) here, especially the convergence after extraction of a subsequence to a limiting space; and they also computed the Hausdorff dimension $2\alpha$ of these limits. This convergence, after extraction of a subsequence, extends to the whole domain of attraction of an $\alpha$-stable law [Mar22, Section 6.4] and it was very recently proved by Curien, Miermont, & Riera [CMR25] that the convergence holds without the need of extracting subsequences.
In the aforementioned bijection, the vertices of a Boltzmann map relate to the negative jumps of the $\nu_{q}$-excursion. From this, one can prove that the number of vertices in such a map conditioned to have $n$ edges concentrates around $\nu_{q}(-1)n$, see once again [Mar22, Section 6.4]. We have considered in [KM23] maps $M^{n,K_{n}}$ with weights in the $\alpha$-stable regime as above and multiconditioned to have $n$ edges and $K_{n}$ vertices. The corresponding excursion $W^{n,K_{n}}$ has the law of a $\nu_{q}$-random excursion conditioned both to have duration $n+1$ and a total $K_{n}-1$ of negative increments. Then Theorem 4.4(i) in [KM23] shows that if we set $r_{n}=(\alpha\Gamma(-\alpha)n)^{1/\alpha}$ and if there exists $\vartheta\in\mathbb{R}$ such that
then
Let us mention that the statement in [KM23] actually applies more generally to non-centred distributions, and the entire domain of attraction of a stable law. The case $\alpha=2$ is also treated there, although in this regime the limit $X^{2,\vartheta,\mathrm{ex}}$ is simply the Brownian excursion independently of $\vartheta$, as it can be checked using Girsanov’s formula.
Corollary 3.1.
and the limit satisfies almost surely:
A motivation of this work, following [KM23], was to make explicit this result, based on the convergence (3.3). In a companion paper, we further study the behaviour of $\mathscr{L}(X^{\alpha,\vartheta,\mathrm{ex}})$ and $M^{\alpha,\vartheta}$ as $\alpha$ and $\vartheta$ vary, partly based on the result in Section 6 below. We prove that the corresponding looptrees interpolate as $\vartheta$ varies from $\infty$- ∞ to $\infty$, or as $\alpha$ varies from $1$ to $2$, between a circle and the Brownian tree, whereas the corresponding maps interpolate between the Brownian tree and the Brownian sphere.
@@ 4. A spinal decomposition and some volume bounds
We then turn to maps. Recall that we let $(M,d_{M},p_{M})$ denote a subsequential limit of discrete random maps, related to a pair $(X^{\mathrm{ex}},Z^{\mathrm{ex}})$, from Theorem 2.1 and Theorem 2.3. In addition, thanks to Skorokhod’s representation theorem, we assume that all convergences in these theorems hold in the almost sure sense. Notice the additional square inside $\psi$ in the next statement compared to the first one, which eventually leads to the factor $1/2$ in the dimensions.
@@
In order to prove these results, we shall first describe in Section 4.1 the path from the root to a uniformly random point in the Lévy looptree, as well as the labels on this path. Then Proposition 4.1 is proved in Section 4.2, whereas Proposition 4.2 is proved in Section 4.3, relying on a technical estimates on Brownian paths stated in Section 4.4 and proved in Section 4.5.
@@ 4.1.   A spinal decomposition
Note that in the case $s\wedge t=s$, or equivalently when $s\prec t$, the first two items are empty and the third one reduces to requiring either $Y_{s}=\inf_{[s,t]}Y$ or $Y_{s-}=\inf_{[s,t]}Y$. It is straightforward to check that these properties hold in each case on the right of the claim, let us prove the direct implication when $Y=X^{\mathrm{ex}}$ by relying on some properties of such a path.
@@
Consider $X$ the unconditioned Lévy process. First almost surely if $\Delta X_{r}>0$, then for every $\varepsilon>0$, we have both $\inf_{[r-\varepsilon,r]}X<X_{r-}$ and $\inf_{[r,r+\varepsilon]}X<X_{r}$. The latter follows from the Markov property and the fact that 00 is regular for $(-\infty,0)$ and the former by invariance under time reversal and the fact that 00 is regular for $(0,\infty)$, see Theorem VII.1 and Corollary VII.5 in [Ber96]. Consequently almost surely if $X$ achieves a local minimum, then it does not jump at this time. Next, the strong Markov property also entails that the local minima are unique. Finally, by [Ber96, Proposition III.2], we know that almost surely every weak record of $X$ obtained by a jump is actually a strict record. Applied to the dual process at a time just after $t$, it shows that almost surely if $X$ realises a local minimum at time $t$ and if $r=\sup\{u<t\colon X_{u}<X_{t}\}$ has $\Delta X_{r}>0$, then $X_{r-}<X_{t}<X_{r}$. All these properties transfer to the bridge $X^{\mathrm{br}}$ by absolute continuity and then to the excursion $X^{\mathrm{ex}}$ by the Vervaat transform. In the rest of this proof, we implicitly assume that they hold true.
Fix $s<t$ such that $d^{a}_{\mathscr{L}(X^{\mathrm{ex}})}(s,t)=0$. We first claim that necessarily $s\prec t$. Indeed, if this fails, then there exists $u\in(s,t)$ such that $X^{\mathrm{ex}}_{u}=\inf_{[s,t]}X^{\mathrm{ex}}<X^{\mathrm{ex}}_{s-}$. Let $r=\sup\{v<u\colon X^{\mathrm{ex}}_{v}<X^{\mathrm{ex}}_{u}\}$. The path $X^{\mathrm{ex}}$ makes a local minimum at time $u$ so $X^{\mathrm{ex}}_{r-}<X^{\mathrm{ex}}_{u}<X^{\mathrm{ex}}_{r}$. By construction we have
$\inf_{[r,s]}X^{\mathrm{ex}}>X^{\mathrm{ex}}_{u}\geq\inf_{[r,t]}X^{\mathrm{ex}}$
and by the properties of $X^{\mathrm{ex}}$, since $r$ is a jump time, then
$X^{\mathrm{ex}}_{r}>\inf_{[r,s]}X^{\mathrm{ex}}.$
Notice that $r=s\wedge t$ so this violates the third requirement of $d^{a}_{\mathscr{L}(X^{\mathrm{ex}})}(s,t)=0$ in the list in the beginning of this proof.
Therefore, we must have $s\prec t$, namely $X^{\mathrm{ex}}_{s-}\leq\inf_{[s,t]}X^{\mathrm{ex}}$. The condition $d^{a}_{\mathscr{L}(X^{\mathrm{ex}})}(s,t)=0$ thus implies $\inf_{[s,t]}X^{\mathrm{ex}}$ be equal to either $X^{\mathrm{ex}}_{s}$ or to $X^{\mathrm{ex}}_{s-}$. By the properties of $X^{\mathrm{ex}}$, if $\Delta X^{\mathrm{ex}}_{s}>0$, then first $\inf_{[s,t]}X^{\mathrm{ex}}<X^{\mathrm{ex}}_{s}$, and second, at time $u=\inf\{r>s\colon X^{\mathrm{ex}}_{r}=X^{\mathrm{ex}}_{s-}\}$, the path $X^{\mathrm{ex}}$ cannot make a local minimum, so we must have $u=t$, which is the first alternative in our claim. In addition, this time $t$ is unique.
@@
Suppose finally that $\Delta X^{\mathrm{ex}}_{s}=0$. Then there is no $r\in(s,t)$ such that $X^{\mathrm{ex}}$ achieves a local minimum at time $r$ with $X^{\mathrm{ex}}_{r}=X^{\mathrm{ex}}_{s}$, nor $X^{\mathrm{ex}}_{s}=X^{\mathrm{ex}}_{r-}<X^{\mathrm{ex}}_{r}$. Therefore if $X^{\mathrm{ex}}_{t-}>X^{\mathrm{ex}}_{s}$, and since $d_{C}(s,t)=0$, then there exists $r\in(s,t)$ such that $\Delta X^{\mathrm{ex}}_{r}>0$ and $X^{\mathrm{ex}}_{s}<X^{\mathrm{ex}}_{r-}$ and $X^{\mathrm{ex}}_{t-}<X^{\mathrm{ex}}_{r}$, which contributes to the second term in (2.6). We conclude that in this case we must have $X^{\mathrm{ex}}_{t-}=X^{\mathrm{ex}}_{s}$. Furthermore, the time $t$ cannot be a jump time, for otherwise $X^{\mathrm{ex}}$ would reach smaller values just before, hence $X^{\mathrm{ex}}_{t}=X^{\mathrm{ex}}_{s}=\inf_{[s,t]}X^{\mathrm{ex}}$. Finally there can only be at most two such times $t$, if $\inf\{t>s\colon X^{\mathrm{ex}}_{t}=X^{\mathrm{ex}}_{s}\}$ is a time of a local minimum since then it is unique at this height. ∎
@@
Let us now describe the “chain of loops” from the root 00 to a given point $u\in[0,1]$ in the looptree, see Figure 6 for a graphical representation. First let $A_{u}$ be the set of pairs of times $(a,b)\in[0,u]\times[u,1]$ such that:
$\text{either}\quad X^{\mathrm{ex}}_{a}=X^{\mathrm{ex}}_{b}=\min_{[a,b]}X^{\mathrm{ex}}\quad\text{or}\quad b=\inf\{t>a\colon X^{\mathrm{ex}}_{t}=X^{\mathrm{ex}}_{a-}\}.$ (4.1)
@@
By Lemma 4.3, pairs $(a,b)\in A_{u}$ are identified in the looptree. These points intuitively correspond to the ancestors of $u$ in the tree. Define then the set of pairs of points that belong to the same cycle on the geodesic from $u$ to the root, with one point on the left and one on the right. Formally, let $B_{u}$ denote the set of pairs of times $(a,b)\in[0,u]\times[u,1]$ for which there exists $s<a$ such that:
$X^{\mathrm{ex}}_{s}>X^{\mathrm{ex}}_{a-}=\min_{[s,a-]}X^{\mathrm{ex}}>X^{\mathrm{ex}}_{b-}=\min_{[s,b-]}X^{\mathrm{ex}}>X^{\mathrm{ex}}_{s-}.$ (4.2)
@@
Pairs in $(a,b)\in B_{u}$ are not identified in the looptree (still by Lemma 4.3). Instead, if $s$ in the time in (4.2) and $t=\inf\{v>s\colon X^{\mathrm{ex}}_{v}=X^{\mathrm{ex}}_{s-}\}>u$, then $(s,t)\in A_{u}$, so they correspond to an ancestor of $u$. This ancestor has a large offspring number $\Delta X^{\mathrm{ex}}_{s}$ which creates a cycle in the looptree. One point on this cycle corresponds to the next ancestor of $u$ and splits the cycle into a left part, to which $a$ belongs, and a right part, to which $b$ belongs. One can easily prove that almost surely, for any $u\in[0,1]$ the set $A_{u}$ is closed and $A_{u}\cup B_{u}$ is the closure of $B_{u}$.
@@
    $u^{-}_{\varepsilon}$ $=\inf\{s<u\colon X^{\mathrm{ex}}_{s-}=X^{\mathrm{ex}}_{s_{\varepsilon}-}=\min_{[s,u]}X^{\mathrm{ex}}\}$ (4.3)
    $u^{+}_{\varepsilon}$ $=\inf\{t>u\colon X^{\mathrm{ex}}_{t}<X^{\mathrm{ex}}_{s-}\}$